Workshop: Representations, cohomology and support spaces: Bielefeld, 29 April–1 May, 2007

Jon F. Carlson (with Eric M. Friedlander):

The exact category of modules of constant Jordan type

Let G be a finite group scheme defined over a field k. The collection of exact sequences of kGmodules which are split at every π-point, provides the category of kG-modules with the structure of an exact category in the sense of Quillen. This structure is a particularly useful setting for studying the realization problem for modules of constant Jordan type. In this the second of two talks on the subject, we concentrate on some specific constructions for modules of constant Jordan type and a couple of cases for which the Grothendieck group of the exact category has been computed.